Esperanza is paid on a daily basis she is paid 312 pesos a day plus commission of 1.5% on all her sales or total sales for the month or 38,000 235 she work for a total of 24 and a half days monthly deductions are four and a half of total monthly paper sss 75 pesos for medicare and 3% of total monthly pay for withholding tax union dues every month is 20 pesos compute for net pay


1
Expert's answer
2021-10-11T01:47:48-0400

Net pay = Total earning - Deductions

Total earning = "24.5*312 + 24.5(1.5\%*38000.235 )=21609.08636"

Deductions ="4.5 \%* 21609.08636+75+3\%* 21609.08636+20 = 1715.681477"

Net pay = Total earning - Deductions "21609.08636-1715.681477=19893.40488"

"=19893 peso"